FOURIER TRANSFORM SPECTRA OF THE A^{2}Π − X^{2}Σ^{+} TRANSITION OF MgBr NEAR 380 NM

\underline{T. HIRAO} and P. F. BERNATH, Centre of Molecular Beam and Laser Chemistry, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

It is well-known that the A^{2}Π state of the magnesium monohalide radicals (MgX, X= F, Cl, Br, I) are predissociated. Spectroscopic results of this predissociation are available, however, only for MgF and MgCl. In this study, the high-resolution emission spectra of the gaseous magnesium bromide (MgBr) radical have been detected. The spectra of MgBr were recorded by using Bruker IFS120HR Fourier transform spectrometer, modified to record double-sided interferograms. The MgBr radical was generated in the microwave discharge of Ar buffer gas (1.5 Torr) with a trace of the vaporized MgBr_{2} salt. The vapor of the MgBr_{2} salt was obtained in a quartz tube heated to about 400^{∘}C with a heating tape. The (0,0) and (1,1) vibrational bands for the A^{2}Π_{1/2} - X^{2}Σ^{+} were clearly identified near 380 nm, but only the (0,0) sub-band for the A^{2}Π_{3/2} - X^{2}Σ^{+} was detected, probably due to a predissociation. The assignment is in progress.